OpenStudy (anonymous):

Let g(x)=x^2(x+4)(x-3) Find g'(x) and g''(x)

OpenStudy (bahrom7893):

I like to make my functions easy before i differentiate them: g(x) = x^2(x^2-3x+4x-12) = x^2(x^2+x-12) = x^4 + x^3 - 12x^2

OpenStudy (bahrom7893):

g'(x) = 4x^3 + 3x^2 - 24x g"(x) = 12x^2 + 6x - 24